OpenStudy (thomasfire55):

Solve 2x 2 + x - 6 = 0 a) {-2, 3/2} b) {-3/2, 2} c) {}

OpenStudy (welshfella):

this function can be factored so the firm (2x + a)(x + b) where a and b are 2 integers can you factor it?

OpenStudy (welshfella):

* to the form , i meant to say!

OpenStudy (welshfella):

have you heard of the ac method of factoring?

OpenStudy (kaylabruhbruh):

wjat are he answers cuz i got x=2

OpenStudy (thomasfire55):

its 2x^2 sorry

OpenStudy (welshfella):

yeah i assumed that ac method is :- multiply first coefficient by the last which is in this case 2 * - 6 = -12 now we need to find what factors of -12 gives us the middle term in the original expression which is +1 ( from the + x so we try 4x - 3x ( = + x0 so we have 2x^2 + x - 6 = 0 2x^2 + 4x - 3x - 6 = 0

OpenStudy (welshfella):

* ( = +x )

OpenStudy (welshfella):

Do you follow that OK?

OpenStudy (thomasfire55):

yea

OpenStudy (welshfella):

OK so now we factor by grouping ( groups of 2) 2x^2 + 4x - 3x - 6 = 0 2(x + 2) = -3(x + 2) = 0

OpenStudy (welshfella):

oh sorry the second line is 2(x + 2) - 3(x + 2) = 0

OpenStudy (welshfella):

what do you notice about what is in each parenthesis)?

OpenStudy (thomasfire55):

their the same

OpenStudy (welshfella):

right so we can take out x + 2 to give (x + 2)(2x - 3) = 0 damn it I made another mistake in the last post its 2x(x + 2) - 3(x + 2) = 0

OpenStudy (welshfella):

so do you know how to find x from (2x - 3(x + 2) = 0

OpenStudy (thomasfire55):

i cant really rember its been awhile sence ive done this stuff

OpenStudy (welshfella):

ok well when to entities are multiplied together and the answer is 0 eother one of them must equal zero so we can split it up into 2x - 3 = 0 and x + 2 = 0 giving x = 3/2 from the first equation and x = what?

OpenStudy (welshfella):

so the answer will be written in the form { ?, 3/2}

OpenStudy (thomasfire55):

so would the awnser be c?

OpenStudy (welshfella):

how can it be c? - when we know that one root is 3/2?

OpenStudy (welshfella):

x + 2 = 0 x = ?

OpenStudy (welshfella):

{} means there are no real roots.

OpenStudy (thomasfire55):

i ment b

OpenStudy (thomasfire55):

a*